# Reward-Robust RLHF in LLMs

### v1

@@ Abstract
As Large Language Models (LLMs) continue to progress toward more advanced forms of intelligence, Reinforcement Learning from Human Feedback (RLHF) is increasingly seen as a key pathway toward achieving Artificial General Intelligence (AGI). However, the reliance on reward-model-based (RM-based) alignment methods introduces significant challenges due to the inherent instability and imperfections of Reward Models (RMs), which can lead to critical issues such as reward hacking and misalignment with human intentions. In this paper, we introduce a reward-robust RLHF framework aimed at addressing these fundamental challenges, paving the way for more reliable and resilient learning in LLMs. Our approach introduces a novel optimization objective that carefully balances performance and robustness by incorporating Bayesian Reward Model Ensembles (BRME) to model the uncertainty set of reward functions. This allows the framework to integrate both nominal performance and minimum reward signals, ensuring more stable learning even with imperfect reward models. Empirical results demonstrate that our framework consistently outperforms traditional RLHF across diverse benchmarks, showing improved accuracy and long-term stability. We also provide a theoretical analysis, demonstrating that reward-robust RLHF approaches the stability of constant reward settings, which proves to be effective in a stochastic-case analysis. Together, these contributions highlight the framework’s potential to enhance both the performance and stability of LLM alignment with RLHF.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal methodology for ensuring the alignment of foundational models with human values and preferences . This technique has proven instrumental in enhancing the capabilities of Large Language Models (LLMs) to generate responses that are more helpful, harmless and honest. The standard RLHF framework is delineated into two principal phases: 1) Reward Model (RM) training from Human/Artificial Intelligence Feedback: In this phase, a RM is trained by preference data, typically utilizing Maximum Likelihood Estimation (MLE). 2) RM-based Reinforcement Learning: Subsequently, the model employs the established deep reinforcement learning algorithm, Proximal Policy Optimization (PPO) Schulman et al. (2017), to refine performance based on the reward function determined in the preceding phase. This structured approach ensures that LLMs operate in a manner that is consistent with ethical guidelines and user expectations, thereby enhancing their capability and trustworthiness in practical applications.
Many researchers believe that the most significant challenges in developing a RM is the inherent biases among annotators—whether human or AI—in the preference data Bai et al. (2022a); Lee et al. (2023). These biases make it exceedingly difficult to create a RM that is fully aligned with the diverse preferences of humanity as a whole. A poor RM may provide incorrect signals for certain data points during the PPO training phase, ultimately compromising the performance of the fine-tuned model. Here, we outline several issues that arise from an imperfect RM: 1) Reward hacking: This occurs when the model exploits loopholes in the reward function, optimizing for unintended behaviors that maximize the reward signal without genuinely improving task performance. 2) Overfitting and underfitting: An overfitted RM can capture noise or specific patterns in the training data that do not generalize well to new data, while an underfitted model fails to capture important patterns altogether, leading to poor decision-making. All the issues mentioned brings us to a critical question:
@@
In this paper, we propose a reward-robust RLHF framework to overcome this challenge. To enhance the model’s robustness to the reward signal while avoiding an overly conservative optimization process, we introduce a novel optimization objective that balances performance and robustness: $J_{\lambda}(\theta)=\lambda J_{\text{perform}}(\theta)+(1-\lambda)J_{\text{robust}}(\theta)$ (Equ.7). Here, $J_{\text{perform}}(\theta)$ measures the nominal performance, based on signals from a nominal reward model (the reward model used as an approximation of the golden one regardless of the uncertainty), while $J_{\text{robust}}(\theta)$ assesses the worst-case performance across all possible reward functions within an uncertainty set. We propose Bayesian Reward Model Ensembles (BRME) to characterize the nominal reward function and the uncertainty set: We train a multi-head RM, where each head outputs the mean and standard deviation of a Gaussian distribution, from which the final reward is sampled. BRME is trained with a Mean Square Error (MSE) loss. BRME has two main advantages over traditional RMs that are trained with MLE loss and output a scalar as the reward: 1) We proved that the standard deviation can reflect the confidence of the head on its output reward, thus the output with the lowest standard deviation can be reasonably chosen as the nominal reward. 2) We show that both the coverage of the reward distribution and the accuracy on preference testset are superior to that with the traditional RMs. Ablation study on the RM training supports the claims above. The proposed reward-robust RLHF framework consistently outperforms standard RLHF across 16 widely-used benchmarks. In long run training processes, the proposed method shows stronger stability and better performance, with an average accuracy of about 4% higher than the traditional method.
@@
Beyond presenting empirical performance results, we also aim to provide deeper theoretical insights. First, we delve into the inherent imperfections of RMs. We argue that, even with an ideal annotator, achieving a perfect reward function is unlikely, and the reward model retrieved is always insufficient. We design a synthetic toy model to demonstrate that, even with an unbiased annotator, it is unlikely that a reward model can be trained to perfectly align with the annotator’s preferences. Second, we provide a theoretical justification for the superiority of our method over standard RLHF. Given that the actual reward is inherently biased—resulting in either over-scoring or under-scoring during training—we show that, in the long run, under-scoring is preferable to over-scoring. Additionally, we conduct a stochastic case analysis within our reward-robust RLHF framework and prove that the resulting policy remains acceptable even in the worst scenarios. Our analysis shows that the robustness regularization term narrows the reward distribution, making the training process closer to the constant reward setting thus more stable, which is preferable to the uncontrollable optimization driven by badly assigned rewards.
The rest of the paper is organized as follows. Section 2 summarizes previous reward-robust research, discussing their influence on our approach and highlighting key differences between their methods and ours. Section 3 introduces the neccessary preliminaries. Section 4 presents synthetic results from a toy model to illustrate the inherent imperfections of reward models. Section 5 formally introduces the proposed reward-robust RLHF framework, detailing the training pipeline of BRME and evaluating its performance across multiple benchmarks. Section 6 provides further discussion, offering insights on critical topics such as the comparison between over-scoring and under-scoring in inherently biased rewards, a stochastic case analysis that explains why the proposed method works, and various ablation studies.
@@ 2 Related Works
2.1 RM-based Alignment in LLMs
@@
The core idea behind RM-based alignment is to use a RM, typically trained on human/AI-annotated data, to guide the optimization of the model’s policy Christiano et al. (2017); Ouyang et al. (2022). Although RM-free algorithms such as DPO Rafailov et al. (2024) and IPO Azar et al. (2024) have also been developed, it is widely recognized that these methods are sub-optimal compared to RM-based approaches Xu et al. (2024); Yan et al. (2024b). While RM-based approach has shown success in several domains, it is not without challenges. One major issue is the potential misalignment between the RM’s scoring output and the true human preferences, which can lead to unintended behaviors, commonly known as reward hacking Amodei et al. (2016). Additionally, the reward model’s generalization ability is often limited by the quality and diversity of the annotated data, which can result in overfitting to specific data points and underfitting to others Lee et al. (2021). These shortcomings underscore the importance of developing better RMs and alternative methods for RM-based RLHF that can better capture the complexity of human preferences and reduce the risk of misalignment.
@@
Recent research has explored various approaches to address these challenges. Shen et al. (2024) introduces a penalty term on the reward, named as contrastive rewards, to improve the effectiveness of RMs. Eisenstein et al. (2023) find that RM ensembles can help mitigate reward hacking in certain scenarios, though they do not establish a universally robust optimization method. Zhang et al. (2024b) proposes a lightweight uncertainty quantification method to assess the reliability of the reward model to avoid over-optimization. Zhang et al. (2024a); Zhai et al. (2023) use low-rank adaptation (LoRA) to increase the diversity of RMs to improve the performance of RLHF.
@@
Our work differs from the above research in several key aspects: 1) We propose a generally robust RLHF framework that outperforms standard RLHF across nearly all benchmarks tested. Within this framework, we employ Bayesian Reward Model Ensembles (BRME) to model an uncertainty set and introduce a trade-off objective that balances nominal performance and robustness, rather than simply using the mean or median of all the rewards as done in Eisenstein et al. (2023); Zhang et al. (2024b). In BRME, we use a Bayesian multi-head RM to characterize the uncertainty of each reward head, which distinguishes our approach from Zhai et al. (2023). The effectiveness of BRME is supported by ablation studies. 2) We demonstrate, using a synthetic toy model, the inherent imperfection of the reward model, even when using an unbiased and ideal annotator, which deepens and supplements previous research on disagreement between annotators Dubey et al. (2024). 3) We provide a theoretical explanation for the superiority of our method through a stochastic-case analysis. Given that the actual reward is inherently biased, tending either to over-score or under-score, we first demonstrate that, over the long term, under-scoring is preferable to over-scoring.
@@ 2.2 Robust Reinforcement Learning
In the realm of robust RL, robustness is primarily focused on addressing uncertainties related to transition, observation, action, or disturbance Moos et al. (2022). Transition-robust approaches deal with uncertainties in system dynamics by deliberately adjusting state transition probabilities Heger (1994); Nilim & El Ghaoui (2005); Satia & Lave Jr (1973); Givan et al. (2000). Observation-robust methods involve distorting the perceived system state to influence policy decisions Zhang et al. (2020); Gleave et al. (2019). Action-robust designs modify system transitions by introducing disturbances to the agent’s actions Tessler et al. (2019). Disturbance-robust strategies account for external forces that introduce uncertainty into system behavior Pinto et al. (2017).
In contrast, there has been comparatively less focus on reward robustness. Xu & Mannor (2006) consider an MDP with an uncertain reward function and then propose a weighted sum between a nominal and a robust performance criterion, which directly inspires our work. The trade-off can be directly made on the expected return Xu & Mannor (2006) or by defining a chance constraint optimization problem Delage & Mannor (2010). Other research targets the rectangularity assumption, which assuming that the uncertainty in each state is independent of all other states, identifying it as a primary source of the reward uncertainty Mannor et al. (2012; 2016); Goyal & Grand-Clement (2023). Vadori et al. (2020) proposes a risk-sensitive RL approach to handle the reward uncertainty by applying Doob decomposition on the reward. Wang et al. (2020) develops a robust RL framework where the agents can only observed perturbed rewards.
In conclusion, due to the focus of past RL applications, particularly in domains like robotics, research on reward robustness has not received sufficient attention. However, in the context of LLM training, the uncertainty inherent in the reward model significantly impacts final performance. Serious issues arising from these shortcomings, such as reward hacking, are significantly hindering the progress of LLM development Chen et al. (2024); Kwon et al. (2023). Therefore, the development of specialized reward-robust algorithms is not only necessary but also urgently required. Our work is a pioneering effort to introduce reward robustness into the LLM RLHF/RLAIF pipeline.
@@ 3 Preliminaries
Standard RLHF. Training LLMs typically involves three stages: Pretraining, SFT, and RLHF. In this section, we outline the standard RLHF-PPO paradigm, widely adopted in advanced research Ziegler et al. (2019); Ouyang et al. (2022).
Beginning with a well-trained SFT model, denoted as $\pi_{0}$, we proceed by sampling two responses from $\pi_{0}$ for each instance in a given prompt set. Subsequently, we compile a set of comparisons $\mathcal{D}=\{(x,a^{+},a^{-})\}$, where $a^{+}$ and $a^{-}$ denote human-preferred and human-dispreferred completions, respectively. The distribution of the preference dataset is assumed to follow the Bradley-Terry model Bradley & Terry (1952), i.e., the probability of response $a^{+}$ is better than $a^{-}$ is given by:
@@
$\max_{\theta}J_{r}(\theta)=\max_{\theta}\sum_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}\left[r(x,a)-\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}\right],$ (4)
where $\pi_{\theta}$ is initialized as $\pi_{0}$ and $\beta$ controls the deviation from the original model. PPO Schulman et al. (2017) is typically used to solve the problem in practice. Algorithms that optimize the policy using a separate reward model are referred to as RM-based alignment.
@@ 4 Inherent Imperfection of Reward Models
The issue of disagreement among human annotators in the RLHF pipeline has been noted in previous research Bai et al. (2022a). More recently, researchers have begun using AI as annotators in what is known as RLAIF, claiming comparable or even superior performance to RLHF Lee et al. (2023); Bai et al. (2022b). However, our evaluation experiments revealed that in the RLHF process, the scoring consistency between human annotators and domain experts was approximately 70%, whereas in the RLAIF process, the consistency between multi-agent AI annotators and domain experts dropped to around 64%. The details can be found in Appendix A.
@@ 5 Reward-robust RLHF
Figure 1: Synthetic experiment results with the toy model. In the standard RLHF pipeline with the green dotted frame, even with a dataset annotated by a global annotator, the obtained RM and the actor trained by PPO stays imperfect. In constrast, in our reward-robust RLHF pipeline with the red dotted frame, with an integration of the nominal reward functions and the uncertainty set, we can obtain the optimal actor within PPO.
@@
The root of the problem in the standard RLHF pipeline lies in the excessive reliance on a single nominal reward model $\widehat{r}$ and the neglect of the uncertainty set (see Equ. 5). Inspired by the previous work in robust RL Xu & Mannor (2006), we introduce a worst-case analysis to eliminate the possibility of disastrous performance and propose a robustness measurement of the policy as (6).
$J_{\text{perform}}(\theta):=\mathbb{E}_{\theta}\{\sum_{h=0}^{N-1}\widehat{r}(s_{h},a_{h})\}$ (5)
$J_{\text{robust}}(\theta):=\min_{r\in\mathcal{R}^{\text{uncertain}}}\mathbb{E}_{\theta}\{\sum_{h=0}^{N-1}r(s_{h},a_{h})\}$ (6)
$J_{\lambda}(\theta):=\lambda J_{\text{perform}}(\theta)+(1-\lambda)J_{\text{robust}}(\theta)$ (7)
In contrast to relying solely on the nominal reward model or the pure worst-case analysis, we use a trade-off term between the performance and the robustness to be our objective function as (7). There are several disadvantages to use (6) as the objective directly. 1) It often leads to an overly conservative solution, resulting in mediocre performance across all situations. 2) The desirability of the solution heavily depends on the precise modeling of the uncertainty set, which is challenging in the context of LLM training. 3) If the nominal reward model $\widehat{r}$ are close to the golden one $r^{*}$, the performance under nominal reward signal can provide valuable insights into predicting performance under $r^{*}$. 4) There is an inherent trade-off between worst-case performance and nominal performance—maximizing one often comes at the expense of the other. On the other hand, by relaxing both criteria, it is possible to achieve a well-balanced solution that offers satisfactory nominal performance while maintaining reasonable robustness to reward uncertainty.
We first demonstrate the effectiveness of the method in the toy model setting. For simplicity, we use different random seeds to train three additional different reward models, forming the uncertainty set $\mathcal{R}^{\text{uncertain}}$, and set $\lambda=0.4$. The results, illustrated in the red dotted frame in Figure 1, show that the optimal actor in PPO is achieved, in contrast to the standard case where only the nominal reward model is used.
In the real LLM setting, the situation becomes more complex. We will delve into the details of uncertainty set modeling, the selection of the nominal reward model, and the evaluation of the quality of the trained reward models in the following subsections. As our framework can be easily integrated into existing RLHF pipelines, we will also provide experimental results to illustrate best practices, such as how to select the trade-off hyperparameter $\lambda$.

### v3

@@ Abstract
As Large Language Models continue to progress toward more advanced forms of intelligence, Reinforcement Learning from Human Feedback is increasingly seen as a key pathway toward achieving Artificial General Intelligence. However, the reliance on reward-model-based alignment methods introduces significant challenges due to the inherent instability and imperfections of Reward Models (RMs), which can lead to critical issues such as reward hacking and misalignment with human intentions. In this paper, we introduce a reward-robust RLHF framework aimed at addressing these fundamental challenges, paving the way for more reliable and resilient learning in LLMs. Our approach introduces a novel optimization objective that carefully balances performance and robustness by incorporating Bayesian Reward Model Ensembles to model the uncertainty set of reward functions. This allows the framework to integrate both nominal performance and minimum reward signals, ensuring more stable learning even with imperfect RMs. Empirical results demonstrate that our framework consistently outperforms baselines across diverse benchmarks, showing improved accuracy and long-term stability. We also provide a theoretical analysis, demonstrating that reward-robust RLHF approaches the stability of constant reward settings, which proves to be acceptable even in a stochastic-case analysis. Together, these contributions highlight the framework’s potential to enhance both the performance and stability of LLM alignment.
@@ 1 Introduction
Reinforcement Learning (RL), particularly in the form of Reinforcement Learning from Human Feedback (RLHF), has become a pivotal methodology for aligning foundational models with human values and preferences. It has played a crucial role in enhancing the capabilities of Large Language Models (LLMs) to generate responses that are more helpful, harmless, and honest, contributing to significant breakthroughs such as OpenAI’s o1 model (OpenAI, 2024). The standard RLHF framework consists of two key phases. First, a Reward Model (RM) is trained on a preference data annotated by Human or Artificial Intelligence (AI) feedback.Following this, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is applied to refine the model’s performance based on the learned reward function. This structured approach ensures that LLMs operate in a manner that is consistent with ethical guidelines and user expectations, thereby enhancing their capability and trustworthiness in practical applications.
The quality of the RM is crucial to the success of PPO. A poor RM may provide incorrect signals for certain data points during the PPO training phase, ultimately compromising the performance of the fine-tuned model. Several issues arise from an imperfect RM. One such issue is reward hacking, where the model exploits flaws in the reward function, optimizing for behaviors that maximize the reward signal without genuinely improving task performance. Another challenge is overfitting and underfitting: an overfitted RM captures noise or specific patterns in the training data that fail to generalize to new data, while an underfitted model may miss important patterns altogether. Additionally, misalignment with human preferences can occur, as biases among annotators—whether human or AI (Bai et al., 2022a; Lee et al., 2023)—make it difficult to align the RM with the diverse preferences of humanity, leading to discrepancies between the model’s behavior and human expectations. All the issues mentioned brings us to a critical question:
@@
In this paper, we propose a reward-robust RLHF framework to overcome this challenge. To enhance the model’s robustness to the reward signal while avoiding an overly conservative optimization process, we introduce a novel objective that strikes a balance between performance and robustness. The performance component is guided by a nominal RM, which serves as an approximation of the ideal reward function. Meanwhile, the robustness component accounts for worst-case scenarios by considering the uncertainty in the reward functions. To capture both the nominal reward and the uncertainty, we introduce Bayesian Reward Model Ensembles (BRME). BRME utilizes a multi-head RM, where each head outputs the mean and standard deviation (std) of a Gaussian distribution, from which the final reward is sampled. BRME has two main advantages over traditional RM that generates a single scalar as the reward: First, we demonstrate that std can reflect the confidence of each head in its output reward, allowing the output with the lowest std to be reasonably selected as the nominal reward. Additionally, we show that BRME outperforms traditional RMs in both the coverage of the reward distribution and accuracy on preference test sets. Ablation study on the RM training supports the claims above. The proposed reward-robust RLHF framework consistently outperforms standard RLHF across several widely-used benchmarks. In long run training processes, the proposed method shows stronger stability and better performance compared with the baselines.
@@
Beyond presenting empirical performance results, we also aim to provide deeper theoretical insights. First, we delve deeper into the inherent imperfections of RMs, arguing that even with an ideal annotator, a perfect reward function is unlikely, and the resulting RM is inherently insufficient. We design a synthetic toy model that directly illustrates these limitations. Second, we provide a theoretical justification for the superiority of our method over standard RLHF. Given that the actual reward is inherently biased—resulting in either over-scoring or under-scoring during training—we show that, in the long run, under-scoring is preferable to over-scoring. Third, we conduct a stochastic case analysis within our reward-robust RLHF framework and prove that the resulting policy remains acceptable, as stochastic scenarios often arise when the RM handles out-of-distribution (OOD) data. Our analysis shows that the robustness regularization term narrows the reward distribution, making the training process closer to the constant reward setting thus more stable, which is preferable to the uncontrollable optimization driven by badly assigned rewards.
Contributions. 1) We propose a reward-robust RLHF framework, introducing BRME to model both nominal rewards and uncertainty, outperforming traditional RMs in reward distribution coverage and accuracy. 2) We provide theoretical insights into RM imperfections, showing through a synthetic toy model that even with an ideal annotator, a perfect reward function is unattainable. 3) We show that under-scoring is preferable to over-scoring in long-term training, given the inherent bias in accessible rewards. Additionally, we conduct a stochastic case analysis, demonstrating that the proposed framework remains effective in OOD scenarios, with the robustness regularization term stabilizing the training process by narrowing the reward distribution.
@@ 2 Related Works
2.1 Reward-Model-based Alignment in LLMs
@@
The core idea behind RM-based alignment is to use a RM, typically trained on human/AI-annotated data, to guide the optimization of the model’s policy (Christiano et al., 2017; Ouyang et al., 2022). Although RM-free algorithms such as DPO (Rafailov et al., 2024) and IPO (Azar et al., 2024) have also been developed, it is widely recognized that these methods are sub-optimal compared to RM-based approaches (Xu et al., 2024; Yan et al., 2024b). While RM-based approach has shown success in several domains, it is not without challenges. One major issue is the potential misalignment between the RM’s output and the true human preferences, which can lead to unintended behaviors, commonly known as reward hacking (Amodei et al., 2016). Additionally, the reward model’s generalization ability is often limited by the quality and diversity of the annotated data, which can result in overfitting to specific data points and underfitting to others (Lee et al., 2021). These shortcomings underscore the importance of developing alternative methods that enhance reward robustness.
@@
Recent research has explored various approaches to address these challenges. Shen et al. (2024) introduces a penalty term on the reward, named as contrastive rewards, to improve the effectiveness of RMs. Eisenstein et al. (2023) finds that RM ensembles can help mitigate reward hacking in certain scenarios. Zhang et al. (2024b) proposes a lightweight uncertainty quantification method to assess the reliability of the reward model to avoid over-optimization. Zhang et al. (2024a); Zhai et al. (2023) uses low-rank adaptation (LoRA) to increase the diversity of RMs to improve the performance of RLHF. Yang et al. (2024b) retains the language model head and add regularization to improve the generalization capability of RMs.
@@
Our work differs from the above research in several key aspects: 1) We propose a generally robust RLHF framework that outperforms standard RLHF across most benchmarks tested. Within this framework, we employ BRME to model an uncertainty set and introduce a trade-off objective that balances nominal performance and robustness, rather than simply using the mean or median of all the rewards as done in Eisenstein et al. (2023); Zhang et al. (2024b). In BRME, we use a Bayesian multi-head RM to characterize the uncertainty of each reward head, which distinguishes from Zhai et al. (2023). The effectiveness of BRME is supported by ablation studies. 2) Using a synthetic toy model, we demonstrate the inherent imperfections of the reward model, even when utilizing an unbiased and ideal annotator. This finding builds upon and complements prior research on annotator disagreement (Dubey et al., 2024). 3) We provide a theoretical explanation for the superiority of our method through a stochastic-case analysis. Given that the actual reward is inherently biased, tending either to over-score or under-score, we first demonstrate that, over the long term, under-scoring is preferable to over-scoring.
@@ 2.2 Robust Reinforcement Learning
In the realm of robust RL, robustness is primarily focused on addressing uncertainties related to transition, observation, action, or disturbance (Moos et al., 2022). Transition-robust approaches deal with uncertainties in system dynamics by deliberately adjusting state transition probabilities (Heger, 1994; Nilim & El Ghaoui, 2005; Satia & Lave Jr, 1973; Givan et al., 2000). Observation-robust methods involve distorting the perceived system state to influence policy decisions (Zhang et al., 2020; Gleave et al., 2019). Action-robust designs modify system transitions by introducing disturbances to the agent’s actions (Tessler et al., 2019). Disturbance-robust strategies account for external forces that introduce uncertainty into system behavior (Pinto et al., 2017).
In contrast, there has been comparatively less focus on reward robustness. Xu & Mannor (2006) consider an MDP with an uncertain reward function and then propose a weighted sum between a nominal and a robust performance criterion, which directly inspires our work. The trade-off can be directly made on the expected return (Xu & Mannor, 2006) or by defining a chance constraint optimization problem (Delage & Mannor, 2010). Other research targets the rectangularity assumption, which assuming that the uncertainty in each state is independent of all other states, identifying it as a primary source of the reward uncertainty (Mannor et al., 2012; 2016; Goyal & Grand-Clement, 2023). Vadori et al. (2020) proposes a risk-sensitive RL approach to handle the reward uncertainty by applying Doob decomposition on the reward. Wang et al. (2020) develops a robust RL framework where the agents can only observed perturbed rewards.
In conclusion, due to the focus of past RL applications, particularly in domains like robotics, research on reward robustness has not received sufficient attention. Notably, in the context of LLM training, the uncertainty inherent in the reward model significantly impacts final performance and hinds the progress of LLM development (Chen et al., 2024; Kwon et al., 2023). Coste et al. (2023) introduce worst-case optimization in LLMs to mitigate overoptimization. However, their approach relies on a purely empirical method to characterize uncertainty through intra-ensemble variance, lacking both structure and interpretability. Therefore, the development of specialized reward-robust algorithms is not only necessary but also urgently required. Our work is one of the pioneering efforts to introduce reward robustness into the LLM RLHF/RLAIF pipeline.
@@ 3 Preliminaries
Standard RLHF. Training LLMs typically involves three stages: Pretraining, Supervised Fine-tuning (SFT), and RLHF. In this section, we outline the standard RLHF-PPO paradigm, widely adopted in advanced research (Ziegler et al., 2019; Ouyang et al., 2022).
Beginning with a well-trained SFT model, denoted as $\pi_{0}$, we proceed by sampling two responses from $\pi_{0}$ for each instance in a given prompt set. Subsequently, we compile a set of comparisons $\mathcal{D}=\{(x,a^{+},a^{-})\}$, where $a^{+}$ and $a^{-}$ denote human-preferred and human-dispreferred completions, respectively. The distribution of the preference dataset is assumed to follow the Bradley-Terry model (Bradley & Terry, 1952), i.e., the probability of response $a^{+}$ is better than $a^{-}$ is given by:
@@
$\max_{\theta}J(\theta)=\max_{\theta}\sum_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}\left[r(x,a)-\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}\right],$ (4)
where $\pi_{\theta}$ is initialized as $\pi_{0}$ and $\beta$ controls the deviation from the original model. PPO (Schulman et al., 2017) is typically used to solve the problem in practice. Algorithms that optimize the policy using a separate reward model are referred to as RM-based alignment.
@@ 4 Inherent Imperfection of Reward Models
The issue of disagreement among human annotators in the RLHF pipeline has been noted in previous research (Bai et al., 2022a). More recently, researchers have begun using AI as annotators in what is known as RLAIF, claiming comparable or even superior performance to RLHF (Lee et al., 2023; Bai et al., 2022b). However, our evaluation experiments revealed that in the RLHF process, the scoring consistency between human annotators and domain experts was approximately 70%, whereas in the RLAIF process, the consistency between multi-agent AI annotators and domain experts dropped to around 64%. The details can be found in Appendix A.
@@ 5 Reward-robust RLHF
Figure 1: Diagram and synthetic experiment results with the toy model. In the standard RLHF pipeline with the upperside gray frame, even with a dataset annotated by a global annotator, the obtained RM and the actor trained by PPO stays imperfect. In constrast, in our reward-robust RLHF pipeline with the downside orange frame, with an integration of the nominal reward functions and the uncertainty set, we can obtain the optimal actor within PPO.
@@
The root of the problem in the standard RLHF pipeline lies in the excessive reliance on a single nominal reward model $\widehat{r}$ and the neglect of the uncertainty set, as shown in Eq. (5). Inspired by the previous work in robust RL (Xu & Mannor, 2006), we introduce a worst-case analysis to eliminate the possibility of disastrous performance and propose a robustness measurement of the policy as Eq. (6).
$J_{\text{perform}}(\theta):=\sum_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}\left\{\widehat{r}(x,a)-\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}\right\},$ (5)
$J_{\text{robust}}(\theta):=\min_{r\in\mathcal{R}^{\text{uncertain}}}\sum_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}\left\{r(x,a)-\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}\right\},$ (6)
$J_{\lambda}(\theta):=\lambda J_{\text{perform}}(\theta)+(1-\lambda)J_{\text{robust}}(\theta).$ (7)
In contrast to relying solely on the nominal reward model or the pure worst-case analysis, we use a trade-off term between the performance and the robustness to be our objective function as Eq. (7). There are several disadvantages to use Eq. (6) as the objective directly. 1) It often leads to an overly conservative solution, resulting in mediocre performance across all situations. 2) The desirability of the solution heavily depends on the precise modeling of the uncertainty set, which is challenging in the context of LLM training. 3) If the nominal reward model $\widehat{r}$ are close to the golden one $r^{*}$, the performance under nominal reward signal can provide valuable insights into predicting performance under $r^{*}$. 4) There is an inherent trade-off between worst-case performance and nominal performance—maximizing one often comes at the expense of the other. On the other hand, by relaxing both criteria, it is possible to achieve a well-balanced solution that offers satisfactory nominal performance while maintaining reasonable robustness to reward uncertainty.
We first demonstrate the effectiveness of the method in the toy model setting. For simplicity, we use different random seeds to train three additional different RMs, forming the uncertainty set $\mathcal{R}^{\text{uncertain}}$, and set $\lambda=0.4$. The results, illustrated in the red dotted frame in Figure 1, show that the optimal actor in PPO is achieved, in contrast to the standard case where only the nominal reward model is used.
In the real LLM scenario, the situation becomes more complex. We will delve into the details of uncertainty set modeling, the selection of the nominal reward function, and the evaluation of BRME and end-to-end PPO performance in the following subsections. As our framework can be easily integrated into existing RLHF pipelines, we will also provide experimental results to illustrate best practices, such as how to select the trade-off hyperparameter $\lambda$.